# Off-equatorial deflections and gravitational lensing. II. In general stationary and axisymmetric spacetimes

### v1

@@ VI Conclusions and Discussions
In this work, we studied the off-equatorial deflections and GL of both null and timelike signals in general SAS spacetimes in the weak deflection limit, with the finite distance effect of source and detector. It is found that as long as the metric functions satisfy certain common separable variable conditions (10), the deflection angles in both the $\phi$ and $\theta$ directions can always be found using the perturbative method. The results as shown in Eqs. (41) and (44), are dual series of $M/r_{0}$ and $r_{0}/r_{s,d}$, and can be directly used in a set of exact GL equations (46). These equations are then solved to find the apparent angles of images in such spacetimes (49).
These results are then applied to the KN, KS, and rotating SV spacetimes to validate the correctness of the method and results, and to find the effect of the spacetime spin as well as that of the characteristic parameter (usually some effective charge) of these spacetimes. It is found that generally both the spacetime spin and charge appear in the second order of both $\Delta\phi$ and $\Delta\theta$, while the non-equatorial effect shows up from the very leading nontrivial order, as illustrated in Eq. (52), (60) and (66).
@@
To show the generality of our method, in Appendix B we supplemented a few other spacetimes whose off-equatorial deflections can be found using our method.
@@ Acknowledgements.
This work is supported by the National Natural Science Fundation of China. The authors thank the helpful discussion with Tingyuan Jiang and Xiaoge Xu.
@@ Appendix B Applications in other SAS spacetimes
A class of spacetimes satisfying these conditions is described by the following line element
@@
$m(r)=\sum_{n=0}^{\infty}\frac{m_{n}}{r^{n}},$ (80)
and it’s seen that for the rotating Bardeen and Tinchev spacetimes, their characteristic parameter appears from the second order of the expansion. While the rotating Ghosh and Hayward ones, they appear from the first and third order respectively.
@@
$-\frac{c_{s}s_{m}^{2}}{2s_{s}^{3}}\left[\left(p_{s}+p_{d}\right)-\frac{2(1+v^{2})m_{0}}{v^{2}r_{0}}\right]^{2}+\mathcal{O}\left(\varepsilon\right)^{3}$ (81b)
@@
$\zeta=8+8v^{2}-12\pi v^{2}-3\pi v^{4}-\frac{2\pi v^{2}(2+v^{2})m_{1}}{m_{0}^{2}}.$ (82)

### v3

@@ VI Conclusions and Discussions
In this work, we studied the off-equatorial deflections and GL of both null and timelike signals in general SAS spacetimes in the weak deflection limit, with the finite distance effect of the source and detector. It is found that as long as the metric functions satisfy certain common separable variable conditions (10), which allows the existence of a GCC, the deflection angles in both the $\phi$ and $\theta$ directions can always be found using the perturbative method. The results, as shown in Eqs. (41) and (44), are dual series of $M/r_{0}$ and $r_{0}/r_{s,d}$, and can be directly used in a set of exact GL equations (46). These equations are then solved to find the apparent angles of images in such spacetimes (49).
These results are then applied to the KN, KS, and RSV spacetimes to validate the correctness of the method and results, and to find the effect of the spacetime spin as well as that of the characteristic parameter (typically an effective charge) of these spacetimes. It is found that generally both the spacetime spin and charge appear in the second order of both $\Delta\phi$ and $\Delta\theta$, while the non-equatorial effect shows up from the very leading nontrivial order, as illustrated in Eqs. (52), (60) and (66).
@@
To show the generality of our method, we supplemented a few other spacetimes whose off-equatorial deflections can be found using our method in Appendix B. We summarize the results computed in the main text and this appendix in Tab. 1 to clearly present the results and the effect of the main parameter(s) in the spacetime on the deflection and/or apparent angles.
Spacetime Metric Eq. Para. Def. angle Eq. Order App. angle Eq. Para. effect Kerr-Newmann (50) $Q$ (52) $2$ (56) $\searrow$ Kerr-Sen (58) $b$ (60) $2$ (63) $\searrow$ Rotating Simpson-Visser (64) $l$ (66) $2$ (69) $\nearrow$ Rotating Bardeen (75) with (76) $g$ (81) with (76) 3 or higher Rotating Hayward (75) with (77) $k$ (81) with (77) 3 or higher Rotating Ghosh (75) with (78) $h$ (81) with (78) $2$ Rotating Tinchev (75) with (79) $j$ (81) with (79) 3 or higher Table 1: Spacetime and their off equatorial deflections. From the second to last columns are the metric equation number, the main parameter of the spacetime, the deflection angles in that spacetime, the lowest order in the deflection angle from which the main parameter appears, and the equation number of the image apparent angles for the spacetime studied in the main text, and lastly the effect of the parameter on the apparent angles, i.e. the monotonicity of the apparent angle as the parameter increases.
The results of this work can in principle be applied to any spacetime with a (non-spherical) axisymmetry. Inside the solar system however, the only known object that can bend the light is the Sun, and yet its dimensionless spin parameter $\hat{a}$ is only of order $10^{-20}$ [59]. Therefore the effect of the spin or the off-equatorial plane effect in the deflection angle and/or apparent angle of images can not be observed for the Sun in the foreseeable future. Instead, most spacetimes studied in this work are black hole spacetimes. Therefore the results are more applicable to more extreme rotating black holes, some particular examples include the M87* [11] and the Sgr A* [12], whose spin parameter $\hat{a}$ could potentially reach a much larger order (roughly a fraction of one). One can also assume that such black holes carry extra parameters such as those appearing in the Kerr-Sen or rotating-Simpson-Visser spacetimes, and attempt to use future observations to constrain the corresponding parameters.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China. The authors thank the helpful discussion with Tingyuan Jiang and Xiaoge Xu. The work of X.Ying is partially supported by the Undergraduate Training Programs for Innovation and Entrepreneurship of Wuhan University.
@@ Appendix B Applications in other SAS spacetimes
The following line element describe a class of spacetimes satisfying these conditions
@@
$m(r)=\sum_{n=0}^{\infty}\frac{m_{n}}{r^{n}}$ (80)
and the coefficients $m_{n}$ for each spacetime can be easily read off from Eqs. (76)-(79). For the rotating Bardeen and Tinchev spacetimes, their characteristic parameter appears from the second order of the expansion. While the rotating Ghosh and Hayward ones appear from the first and third order respectively.
@@
$-\frac{c_{s}s_{m}^{2}}{2s_{s}^{3}}\left[\left(p_{s}+p_{d}\right)-\frac{2(1+v^{2})m_{0}}{v^{2}r_{0}}\right]^{2}+\mathcal{O}\left(\varepsilon\right)^{3},$ (81b)
@@
$\zeta=8+8v^{2}-12\pi v^{2}-3\pi v^{4}-\frac{2\pi v^{2}(2+v^{2})m_{1}}{m_{0}^{2}}$ (82)
and all $m_{i}$ should be read off from corresponding spacetime.
@@
For future reference, we also test the applicability of our methodology to SAS but non-asymptotically flat spacetimes, such as the KN-(anti)de Sitter spacetime described by [60]
$\mathrm{d}s^{2}=$ $\frac{\Delta_{\theta}}{\rho^{2}\Xi^{2}}\left[a\mathrm{d}t-\left(r^{2}+a^{2}\right)\mathrm{d}\phi\right]^{2}-\frac{\Delta_{r}}{\rho^{2}\Xi^{2}}\left(\mathrm{d}t-a\sin^{2}\theta\mathrm{d}\phi\right)^{2}{}$
$+\rho^{2}\left(\frac{\mathrm{d}r^{2}}{\Delta_{r}}+\frac{\mathrm{d}\theta^{2}}{\Delta_{\theta}}\right),$ (85)
where
$\rho^{2}=r^{2}+a^{2}\cos^{2}\theta,\,\Delta_{\theta}=1+\frac{\Lambda}{3}a^{2}\cos^{2}\theta,{}$
$\Delta_{r}=\left(r^{2}+a^{2}\right)\left(1-\frac{\Lambda r^{2}}{3}\right)-2Mr+Q^{2},\,\Xi=1+\frac{\Lambda}{3}a^{2},{}$
and Kerr-Taub-NUT spacetime with metric [61]
$\mathrm{d}s^{2}=$ $-\frac{\Delta-a^{2}\sin^{2}\theta}{\Sigma}\mathrm{d}t^{2}+\frac{2[\Delta\chi-a(\Sigma+a\chi)\sin^{2}\theta]}{\Sigma}\mathrm{d}t\mathrm{d}\phi{}$
$+\frac{(\Sigma+a\chi)^{2}\sin^{2}\theta-\chi^{2}\Delta}{\Sigma}\mathrm{d}\phi^{2}+\frac{\Sigma}{\Delta}\mathrm{d}r^{2}+\Sigma\mathrm{d}\theta^{2},$ (86)
where
$\Sigma=r^{2}+(\hat{n}+a\cos\theta)^{2},{}$
$\Delta=r^{2}-2Mr+a^{2}-\hat{n}^{2},{}$
$\chi=a\sin^{2}\theta-2\hat{n}\cos\theta.{}$
It is found that they also satisfy the separation requirements (10) and therefore the deflection of both null and timelike rays in the equatorial or off-equatorial plane in them can be treated using our method. Lastly for C-type metrics which does not have the reflective symmetry about the equatorial plane, including the KN-(A)dS C-metric [62]
$\mathrm{d}s^{2}=$ $\frac{1}{H^{2}}\Bigg{\{}-\frac{f(r)}{\Sigma}\left(\frac{\mathrm{d}t}{\alpha}-a\sin^{2}\theta\frac{\mathrm{d}\phi}{K}\right)^{2}+\frac{\Sigma}{f(r)}\mathrm{d}r^{2}{}$
$+\frac{\Sigma r^{2}}{h(\theta)}\mathrm{d}\theta^{2}+\frac{h(\theta)\sin^{2}\theta}{\Sigma r^{2}}\left[\frac{a\mathrm{d}t}{\alpha}-\left(r^{2}+a^{2}\right)\frac{\mathrm{d}\phi}{K}\right]^{2}\Bigg{\}},$ (87)
where
$f(r)=\left(1-A^{2}r^{2}\right)\left(1-\frac{2m}{r}+\frac{a^{2}+e^{2}}{r^{2}}\right)+\frac{r^{2}+a^{2}}{l^{2}},{}$
$h(\theta)=1+2mA\cos\theta+\left[A^{2}\left(a^{2}+e^{2}\right)-\frac{a^{2}}{l^{2}}\right]\cos^{2}\theta,{}$
$\Sigma=1+\frac{a^{2}}{r^{2}}\cos^{2}\theta,\,H=1+Ar\cos\theta,{}$
and its subcases with $A\neq 0$, we found that only for null but not the timelike rays, the separation requirements (10) can be met and therefore the deflection can be studied using our approach. For these metrics, however, we will not list the formulas in the $\theta$ and $\phi$ directions until more valuable applications arise.